OpenStudy (anonymous):

Write a quadratic variation equation if g(x) varies directly with x2, and g(x) = 108 when x = 6. If possible, please explain so I can learn.

OpenStudy (mathmale):

This is direct variation; this means that if x increases, so does y. The key is to translate that problem into appropriate formulas. The general form of the equation for "quadratic variation" is y=k*x^2, where k is just a constant and is called the "constant of proportionality." Here, we'll call this function g(x) = k*x^2. We are told that g(6) = 108. Substitute 108 for g(x) in the above general equation and substitute 6 for x. Then solve the resulting equation for k.